OpenStudy (anonymous):

Write x2 - 6x + 7 = 0 in the form (x - a)2 = b, where a and b are integers.

OpenStudy (zzr0ck3r):

for \(x^2+bx+c=0\) we have \((x-\frac{-b}{2})=(\frac{b}{2})^2-c\)

OpenStudy (anonymous):

?

OpenStudy (zzr0ck3r):

do you see that if you know what b and c are you can use that form to get the answer?

OpenStudy (anonymous):

oh. thx

OpenStudy (zzr0ck3r):

so for you b = -6, and c = 7

OpenStudy (anonymous):

sorry, i don't get how i solve with the x

OpenStudy (anonymous):

hold on...

OpenStudy (anonymous):

i got 21... i did it wrong

OpenStudy (zzr0ck3r):

\((x-\frac{-(-6)}{2})^2=(\frac{6}{2})^2-7\)

OpenStudy (zzr0ck3r):

\((x-3)^2=2\)

OpenStudy (zzr0ck3r):

you are not trying to get a number....reread the question.

OpenStudy (anonymous):

oh. no, u got it. thank you so much :)

OpenStudy (anonymous):

i get it now. thx